A Farm is divided into 3 regions. The area of one region is twice as large as the area of the smallest region, and the area of the third region is 4 acres less than 3 times the area of the smallest region. Total acreage is 512 acres. Find the area of each of the 3 regions.
Answer by Paul(988)

Let the smallest region be x
Let the one region be 2x
LEt the third region be 3x-4
Total = 512
x+2x+3x-4=512
6x-4=512
6x=516
x=86

2(86)=172
3(86)-4=254

Hence, the smallest region is 86 acres, the second region is 172 acres and the third region is 254 acres.
PAul.